# Possible explanations of the observed $\Lambda_{c}$ resonances

### v1

@@ III.4 $1D$-wave $\rho$-mode excitations
$\Gamma[\Lambda_{c}\omega]$ 31.25
$\Gamma[\Xi_{c}^{\prime}K]$ 0.02
@@
Then accounting for the uncertainty of the predicted masses, which may bring uncertainties to the theoretical results, we plot the decay properties of the $1D$-wave $\rho$-mode $\Lambda_{c}$ as functions of masses within the range of $M=(3.00-3.15)$ GeV, as shown in Fig. 3. We can find that the total decay widths vary within the scope of $\Gamma_{\mathrm{Total}}<200$ MeV as the mass increasing. In addition, when the masses of the $1D$-wave $\rho$-mode $\Lambda_{c}$ states are above the threshold of $\Lambda_{c}\omega$, the partial decay width of this channel will be sizable and increases dramatically with the mass.
@@ III.5 $2S$-wave excitations
$\Gamma[\Sigma_{c}|J^{P}=1/2^{-},0\rangle_{\lambda}\pi]$ 13.15
$\Gamma[\Sigma_{c}|J^{P}=1/2^{-},1\rangle_{\lambda}\pi]$ 0.00
$\Gamma[\Sigma_{c}|J^{P}=3/2^{-},1\rangle_{\lambda}\pi]$ 0.00
$\Gamma[\Sigma_{c}|J^{P}=3/2^{-},2\rangle_{\lambda}\pi]$ 0.01
@@ III.6 $2P$-wave $\lambda$-mode excitations
From the table, it is known that the predicted masses of the two $2P$-wave $\lambda$-mode $\Lambda_{c}$ baryons are about $M\sim 3.00$ GeV, which is close to the measured mass of the observed state $\Lambda_{c}(2940)^{+}$. As the possible assignment, it is crucial to investigate the decay behaviors of the two $2P$-wave $\lambda$-mode $\Lambda_{c}$ baryons. Considering the uncertainties of the predicted masses of $\Lambda_{c1}|J^{P}=1/2^{-},1\rangle_{\lambda}$ and $\Lambda_{c1}|J^{P}=3/2^{-},1\rangle_{\lambda}$, we plot the decay width as a function of the mass in the range of $M=(2.90-3.05)$ GeV in Fig. 5.
@@
In conclusion, our calculation indicates that the strong decay properties of $\Lambda_{c1}|J^{P}=3/2^{-},1\rangle_{\lambda}$ is in good agreement with the nature of $\Lambda_{c}(2940)$. Hence, $\Lambda_{c}(2940)$ can be explained as $\Lambda_{c1}|J^{P}=3/2^{-},1\rangle_{\lambda}$.
@@ III.7 $2P$-wave $\rho$-mode excitations
The decay width of the state $\Lambda_{c1}|J^{P}=3/2^{-},2\rangle_{\rho}$ is about $\Gamma\simeq 293$ MeV. Its strong decays are governed by the $\Sigma_{c}\rho$ channel with the branching fraction around $\sim 60\%$. Meanwhile, this state has a sizable decay rate decaying into the $\Sigma_{c}|J^{P}=1/2^{-},1\rangle_{\rho}\pi$ channel, and the predicted branching ratio is about $\sim 12\%$. However, the $\bar{\Lambda}_{c}|J^{P}=3/2^{-},2\rangle_{\rho}$ might be too broad to observed in experiments.
@@
Similarly, we also plot the partial decay widths of the $2P$-wave $\rho$-mode $\Lambda_{c}$ baryons as a function of the mass in region of $M=(3.10-3.30)$ GeV. The sensitivities of the decay properties if these states to their masses are shown in Fig. 6. From the figure, the partial widths of dominant decay channels for most of the states vary gently with mass increasing. In addition, if the $2P$-wave $\rho$-mode $\Lambda_{c}$ baryons are above the threshold of $\Sigma_{c}\rho$, most of the states can decay via $\Sigma_{c}\rho$ with a partial width about several MeV or more.
@@ IV Summary
In the present work, we have systematically studied the strong decay properties of the low-lying $1P$-, $1D$-, $2S$- and $2P$-wave $\Lambda_{c}$ baryons in the framework of the quark pair creation model within the $j-j$ coupling scheme. Our main results are summarized as follows.
@@
For the $1P$-wave $\rho$-mode $\Lambda_{c}$ baryons, $\Lambda_{c}|J^{P}=1/2^{-},0\rangle_{\rho}$ is most likely to be a very narrow state, and the $\Sigma_{c}\pi$ decay mode almost saturates its total decay widths. Hence, the $\Lambda_{c}\pi\pi$ final state may be a ideal decay channel to explore $\Lambda_{c}|J^{P}=1/2^{-},0\rangle_{\rho}$ in future experiments. The states $\Lambda_{c}|J^{P}=1/2^{-},1\rangle_{\rho}$ and $\Lambda_{c}|J^{P}=3/2^{-},1\rangle_{\rho}$ are probably two quite broad states with a width of around $\Gamma_{\mathrm{Total}}\sim 780$ MeV. Their dominant decay channels are $\Sigma_{c}\pi$ and $\Sigma_{c}^{*}\pi$, respectively. Considering the decay widths too broad, the two states may be difficult to be observed in experiments. The total decay widths of $\Lambda_{c}|J^{P}=3/2^{-},2\rangle_{\rho}$ and $\Lambda_{c}|J^{P}=5/2^{-},2\rangle_{\rho}$ are several tens MeV, and their strong decays are governed by the $\Sigma_{c}\pi$ and $\Sigma_{c}^{*}\pi$ channels. Combining the mass and the decay properties of the newly observed state $\Lambda_{c}(2910)^{+}$, both $\Lambda_{c}|J^{P}=3/2^{-},2\rangle_{\rho}$ and $\Lambda_{c}|J^{P}=5/2^{-},2\rangle_{\rho}$ are good candidates. Further determine which one of the two is it, the partial decay width ratio between $\Sigma_{c}\pi$ and $\Sigma_{c}^{*}\pi$ may be a good criterion.

### v2

@@ III.4 $1D$-wave $\rho$-mode excitations
$\Gamma[\Lambda_{c}\omega]$ - 31.25
$\Gamma[\Xi_{c}^{\prime}K]$ - 0.02
@@
Then accounting for the uncertainty of the predicted masses, which may bring uncertainties to the theoretical results, we plot the decay properties of the $1D$-wave $\rho$-mode $\Lambda_{c}$ as functions of masses within the range of M=(3.003.15)M=(3.00-3.15) GeV, as shown in Fig. 3. We can find that the total decay widths vary within the scope of $\Gamma_{\mathrm{Total}}<200$ MeV as the mass increasing. In addition, when the masses of the $1D$-wave $\rho$-mode $\Lambda_{c}$ states are above the threshold of $\Lambda_{c}\omega$, the partial decay width of this channel will be sizable and increases dramatically with the mass.
@@ III.5 $2S$-wave excitations
$\Gamma[\Sigma_{c}|J^{P}=1/2^{-},0\rangle_{\lambda}\pi]$ - - 13.15
$\Gamma[\Sigma_{c}|J^{P}=1/2^{-},1\rangle_{\lambda}\pi]$ - - 0.00
$\Gamma[\Sigma_{c}|J^{P}=3/2^{-},1\rangle_{\lambda}\pi]$ - - 0.00
$\Gamma[\Sigma_{c}|J^{P}=3/2^{-},2\rangle_{\lambda}\pi]$ - - 0.01
@@ III.6 $2P$-wave $\lambda$-mode excitations
From the table, it is known that the predicted masses of the two $2P$-wave $\lambda$-mode $\Lambda_{c}$ baryons are about $M\sim 3.00$ GeV, which is close to the measured mass of the observed state $\Lambda_{c}(2940)^{+}$. As the possible assignment, it is crucial to investigate the decay behaviors of the two $2P$-wave $\lambda$-mode $\Lambda_{c}$ baryons. Considering the uncertainties of the predicted masses of $\Lambda_{c1}|J^{P}=1/2^{-},1\rangle_{\lambda}$ and $\Lambda_{c1}|J^{P}=3/2^{-},1\rangle_{\lambda}$, we plot the decay width as a function of the mass in the range of M=(2.903.05)M=(2.90-3.05) GeV in Fig. 5.
@@
In conclusion, our calculation indicates that the strong decay properties of $\Lambda_{c1}|J^{P}=3/2^{-},1\rangle_{\lambda}$ is in good agreement with the nature of $\Lambda_{c}(2940)$, and $\Lambda_{c1}|J^{P}=3/2^{-},1\rangle_{\lambda}$ can be a good candidate. It should be pointed out that the threshold of the main decay channel $DN$ is close to the mass of $\Lambda_{c}(2940)$, which may suggest the importance of the coupled-channel effects for understanding the $\Lambda_{c}(2940)^{+}$ state.
@@ III.7 $2P$-wave $\rho$-mode excitations
The decay width of the state $\Lambda_{c1}|J^{P}=3/2^{-},2\rangle_{\rho}$ is about $\Gamma\simeq 293$ MeV. Its strong decays are governed by the $\Sigma_{c}\rho$ channel with the branching fraction around $\sim 60\%$. Meanwhile, this state has a sizable decay rate decaying into the $\Sigma_{c}|J^{P}=1/2^{-},1\rangle_{\rho}\pi$ channel, and the predicted branching ratio is about $\sim 12\%$. However, the $\Lambda_{c1}|J^{P}=3/2^{-},2\rangle_{\rho}$ might be too broad to observed in experiments.
@@
Similarly, we also plot the partial decay widths of the $2P$-wave $\rho$-mode $\Lambda_{c}$ baryons as a function of the mass in region of M=(3.103.30)M=(3.10-3.30) GeV. The sensitivities of the decay properties if these states to their masses are shown in Fig. 6. From the figure, the partial widths of dominant decay channels for most of the states vary gently with mass increasing. In addition, if the $2P$-wave $\rho$-mode $\Lambda_{c}$ baryons are above the threshold of $\Sigma_{c}\rho$, most of the states can decay via $\Sigma_{c}\rho$ with a partial width about several MeV or more.
@@ IV Summary
In the present work, we have systematically studied the strong decay properties of the low-lying $1P$-, $1D$-, $2S$- and $2P$-wave $\Lambda_{c}$ baryons in the framework of the quark pair creation model within the jjj-j coupling scheme. Our main results are summarized as follows.
@@
For the $1P$-wave $\rho$-mode $\Lambda_{c}$ baryons, $\Lambda_{c}|J^{P}=1/2^{-},0\rangle_{\rho}$ is most likely to be a very narrow state, and the $\Sigma_{c}\pi$ decay mode almost saturates its total decay widths. Hence, the $\Lambda_{c}\pi\pi$ final state may be a ideal decay channel to explore $\Lambda_{c}|J^{P}=1/2^{-},0\rangle_{\rho}$ in future experiments. The states $\Lambda_{c}|J^{P}=1/2^{-},1\rangle_{\rho}$ and $\Lambda_{c}|J^{P}=3/2^{-},1\rangle_{\rho}$ are probably two quite broad states with a width of around $\Gamma_{\mathrm{Total}}\sim 780$ MeV. Their dominant decay channels are $\Sigma_{c}\pi$ and $\Sigma_{c}^{*}\pi$, respectively. Considering the decay widths too broad, the two states may be difficult to be observed in experiments. The total decay widths of $\Lambda_{c}|J^{P}=3/2^{-},2\rangle_{\rho}$ and $\Lambda_{c}|J^{P}=5/2^{-},2\rangle_{\rho}$ are several tens MeV, and their strong decays are governed by the $\Sigma_{c}\pi$ and $\Sigma_{c}^{*}\pi$ channels. Combining the mass and the decay properties of the newly observed state $\Lambda_{c}(2910)^{+}$, both $\Lambda_{c}|J^{P}=3/2^{-},2\rangle_{\rho}$ and $\Lambda_{c}|J^{P}=5/2^{-},2\rangle_{\rho}$ may be good candidates. Further determine which one of the two is it, the partial decay width ratio between $\Sigma_{c}\pi$ and $\Sigma_{c}^{*}\pi$ may be a good criterion.